Two stocks, A and B, are perfectly negatively correlated. The variance of the stocks A and B are .3 and .4 respectively. What proportion should you invest in A to minimize the risk of your portfolio?
Answer #1

sda is standard deviation of Stock A
sdb is standard deviation of Stock B
w is weight of stock A
For minimizing the risk of the portfolio:
w*sda-(1-w)*sdb=0
=>w=sdb/(sda+sdb)
=>Weight of Stock A=sqrt(0.4)/(sqrt(0.3)+sqrt(0.4))=0.535898385